# Physics-Aware Semi-Supervised Underwater Image Enhancement

### v2

@@ 3.2 IFM-Inspired Semi-Supervised Learning Framework
To overcome the challenge of insufficient real-world underwater images, we adopt a semi-supervised learning framework (see Fig.3), inspired by the IFM theory. This framework includes a bi-directional supervised scheme and an unsupervised scheme. The bi-directional scheme contains a forward-enhancement module, which learns from reference images and uses the estimated IFM parameters to perform UIE, and a backward-degradation module, which uses estimated parameters to degrade the reference images to their real-world degraded counterparts. As a result, the bi-directional supervised scheme can learn a more accurate estimation of the parameters using limited real-world data than that learned using a supervised method with a large number of synthetic data. Also, the generalization of the model trained is stronger than that trained using a supervised method with limited real-world data. On the other hand, the unsupervised scheme exploits the unlabeled real-world underwater images that we collect. A second set of images with different degrees of degradation are generated from these images based on the IFM. Two sets of parameters can be estimated using both sets of images respectively. Each set is utilized as the reference for the other set because their IFMs are related.
@@
Figure 3: The proposed semi-supervised learning framework, which comprises a bi-directional supervised learning scheme and an unsupervised learning scheme, used for training our PA-UIENet.
@@
Due to the ill-posed nature of the IFM, however, only $\mathcal{L}_{fwd}$ may be insufficient for ensuring the validity of the estimated parameters. Therefore, we propose a backward-degradation learning module which degrades clear reference images using the estimated parameters to their degraded counterparts $\hat{I}^{c}(x)$. The supervision over the backward-degradation is achieved by minimizing the loss function:
@@
Unsupervised Learning. To leverage unlabeled underwater images, generative networks [21, 15] were used. However, they were difficult to train and produced unstable results. Since unsupervised learning methods [23, 24] are able to directly learn from degraded images without requiring reference images, we are motivated to adopt an unsupervised learning scheme on top of the IFM.
Given an unlabeled real-world degraded image $I^{c}_{1}(x)$, the IFM can be expressed as:
$I^{c}_{1}(x)=J^{c}_{1}(x)t^{c}_{1}(x)+(1-t^{c}_{1}(x))A^{c}_{1}.$ (6)
As performed in [23], a new degraded image $I^{c}_{2}(x)$ can be purposely derived according to:
$I^{c}_{2}(x)=\alpha I^{c}_{1}(x)+(1-\alpha)A^{c}_{1},$ (7)
@@
$I^{c}_{2}(x)=J^{c}_{1}(x)(\alpha t^{c}_{1}(x))+(1-\alpha t^{c}_{1}(x))A^{c}_{1}.$ (8)
Eq. (8) also satisfies the underwater IFM but presents the more severe degradation than the $I^{c}_{1}(x)$, generated from the same underlying clear image $J^{c}_{1}(x)$, because $\alpha t^{c}_{1}(x)<t^{c}_{1}(x)$ always holds true. Therefore, the parameters $\hat{t}^{c}_{2}(x)$ and $\hat{A}^{c}_{2}$ estimated using the PA-UIENet from image $I^{c}_{2}(x)$ will be close to $\alpha\hat{t}^{c}_{1}(x)$ and $\hat{A}^{c}_{1}$ respectively, where $\hat{t}^{c}_{1}(x)$ and $\hat{A}^{c}_{1}$ are the parameters estimated from $I^{c}_{1}(x)$ [23]. Correspondingly, two loss functions are defined as:
$\mathcal{L}_{T}=||\hat{t}^{c}_{2}(x)-\alpha\hat{t}^{c}_{1}(x)||_{2}^{2},$ (9)
$\mathcal{L}_{A\text{-}unsup}=||\hat{A}^{c}_{2}-\hat{A}^{c}_{1}||_{2}^{2}.$ (10)
@@ 4 Experimental Settings
Baselines. We compared the proposed method with four prior-based approaches which were developed on top of the IFM, including DCP [2], UDCP [3], GDCP [4] and DHL [8], two prior-based methods which did not consider the IFM, including Histogram Equalization (HE) and Retinex [11], and two supervised deep learning approaches, including WaterNet [16] and UColor [17].
@@
Data Sets. We conducted a series of experiments on five testing sets derived from four publicly available real-world underwater data sets: UIEB [16], SUIM-E [18], RUIE [43] and SQUID [8]. The UIEB data set originally contains 890 pairs of degraded images, in which each reference image was selected from the results produced by applying 12 algorithms to a degraded image. Ten pairs of UIEB [16] images were discarded because the degraded images in these pairs had been included in the rest pairs. Only 880 pairs of UIEB images were used in this study. The SUIM-E data set comprises 1,525 pairs of degraded underwater images and a testing set of 110 labeled images. Three subsets are comprised of the RUIE data set, including UIQS, UCCS and UHTS, which include 3,630, 300 and 300 unlabeled images, respectively. The SQUID data set comprises 114 underwater images captured using a stereo camera, divided into the Michmoret, Katzaa, Nachsholim and Satil subsets.
@@
Regarding the supervised learning, we randomly selected 720 pairs from the 880 pairs UIEB [16] images and used these as the training set. For the unsupervised learning, we collected 1,934 unlabeled real-world underwater images, which cover diverse underwater scenes. In total, our training set contained 2,654 images. Among the rest of the UIEB [16] data set, 80 pairs of images were randomly selected as the validation set, while the remaining pairs were used as the testing set which is referred to as Test-U80. The UIEB data set also consists of 60 challenging images without reference, which were used as the second testing set, referred to as Test-C60. A third testing set was obtained from the testing set of the SUIM-E data set [18], which contains 110 pairs of labeled underwater images and is named as Test-S110. For the RUIE [43] data set, the UCCS was utilized as the fourth testing set, denoted as Test-UCCS. Besides, 53 SQUID [8] images captured using the camera at the right hand side were randomly selected and were used as the fifth testing set, namely, Test-R53.
@@
Evaluation Metrics. For the labeled testing data, we used both Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [44] to assess the quality of an enhanced image with regard to the reference image. Regarding the unlabeled testing data, we used the Underwater Image Quality Measure (UIQM) [45] and Underwater Color Image Quality Evaluation (UCIQE) [46] metrics. When the SQUID [8] data set was tested, the Average Reproduction Angular Error ($\bar{\psi}$) [8] was used to evaluate the quality of color restoration. Since transmission maps are related to the scene distance, we used the Pearson Correlation Coefficient (PCC) calculated between the estimated transmission maps and the ground-truth depth map to assess the performance of transmission estimation [8].
@@
Implementation Details. We implemented our method using Pytorch and conducted experiments on Ubuntu 20.04 with a GeForce RTX 3080 graphics processing unit. During the training process, the images were resized to the resolution of $256\times 256$ pixels. The number of filters in the RCT was 16. For the $Enc_{1}$ to $Dec_{1}$ in the T-stream, the numbers of filters were 64, 128, 256, 512, 512, 256, 128, 64 and 64 in turn. Regarding the A-stream, the dimension of each token was 384 and the number of attention heads in the MHSA was 6. We trained the PA-UIENet using the AdamW [47] optimizer. The learning rate and batch size were set to 1e-4 and 6 respectively. Four weighting factors, including $\lambda_{1},\lambda_{2},\lambda_{3}$ and $\lambda_{unsup}$ were set to 0.001, 0.005, 10 and 0.001 respectively. We first trained the PA-UIENet using the bi-directional supervised scheme for 3,000 iterations as a warm-up stage. Then the semi-supervised scheme was used to train it for 150,000 iterations. When the semi-supervised learning was carried out, in particular, the unsupervised scheme was performed for 30 iterations after the bi-directional supervised scheme was conducted for every 120 iterations.
5 Experimental Results
@@
In this section, we will report the results obtained in the underwater image enhancement evaluation, parameter estimation and color restoration evaluation and ablation study.
5.1 Evaluation on Underwater Image Enhancement
The quantitative evaluation of eight baselines and the proposed method was conducted for the UIE task. The results are reported in Tables 1. As can be seen, our method outperformed both six prior-based methods and two supervised deep learning approaches on two labeled data sets, including Test-U80 and Test-S110, when the PSNR and SSIM metrics were used. On the other hand, the UIQM and UCIQE metrics were also used in the evaluation experiment where we did not utilize reference images. In this case, the three deep learning methods were inferior to the prior-based approaches. This finding is consistent with those reported in the existing studies [16, 18]. Specifically, the performance of our method was normally comparable to that of the two deep learning approaches, i.e., WaterNet [16] and UColor [17], and was comparable to that of the prior-based methods. As demonstrated in [16, 18, 25], both the UIQM and UCIQE metrics are not always consistent with the quality of the images processed by UIE methods. It should be noted that, however, the prior-based methods did not produce stable results across the five testing sets. This should be attributed to the fact that these methods normally struggle with adapting to diverse underwater scenes or estimating valid parameters when the scene is not suited to the prior assumption [4, 16].
@@ 5.1 Evaluation on Underwater Image Enhancement
Table 1: The quantitative evaluation of eight baselines and our method on five real-world testing sets. The best and second best results are highlighted in the Red Bold and Blue Italic fonts respectively.
Method Test-U80 Test-S110 Test-C60 Test-UCCS Test-R53
@@
DCP 12.48 63.34 1.35 0.51 14.52 67.38 1.36 0.57 0.56 0.51 0.03 0.44 -2.05 0.42
UDCP 9.37 33.62 2.69 0.59 10.04 34.28 1.23 0.60 1.24 0.54 1.83 0.57 0.76 0.59
GDCP 12.46 67.41 2.17 0.56 14.67 71.29 2.12 0.64 1.49 0.60 1.24 0.56 5.50 0.50
DHL 14.49 62.81 3.19 0.64 14.96 69.68 2.35 0.66 1.85 0.64 2.98 0.63 0.22 0.64
HE 16.48 77.77 3.58 0.67 15.37 74.76 2.97 0.68 2.32 0.65 3.93 0.67 1.71 0.64
Retinex 1038 61.25 2.33 0.48 8.29 53.44 0.85 0.45 1.21 0.46 2.03 0.47 -2.03 0.51
WaterNet 17.02 70.27 2.11 0.50 18.80 75.19 1.47 0.54 1.28 0.51 1.25 0.45 0.09 0.45
UColor 20.93 85.19 3.44 0.60 20.23 84.23 2.30 0.63 2.08 0.58 3.22 0.55 1.28 0.54
Ours 22.82 88.87 3.60 0.57 20.46 87.22 2.35 0.60 2.02 0.56 3.24 0.53 1.92 0.56
The qualitative comparison of the results produced by the eight baselines and our method is shown in Fig.4. It can be seen that the enhanced images derived using our method presented the better visual quality, compared with the images enhanced by the baselines. Our method not only effectively removed the color casts in the degraded images but also greatly increased the visibility of them. In contrast, the prior-based methods usually introduced artifacts that decreased the visual quality while the learning-based methods could not eliminate the color casts completely.
Figure 4: The results obtained using eight baselines and our method on five degraded images in Test-U80, Test-S110, Test-C60, Test-UCCS and Test-R53, respectively, shown from top to bottom.
5.2 Evaluation on Transmission Estimation and Color Restoration

### v3

@@ III-C IFM-Inspired Semi-supervised Learning Framework
To overcome the challenge of insufficient real-world underwater images, we adopt a semi-supervised learning framework (see Fig. 3), inspired by the revised IFM [2] theory. This framework includes a bi-directional supervised scheme and an unsupervised scheme. The bi-directional scheme contains a forward-enhancement module, which learns from reference images and uses the estimated parameters of the revised IFM to perform UIE, and a backward-degradation module, which uses the estimated parameters to degrade the reference images to their real-world degraded counterparts. As a result, the bi-directional supervised scheme can learn a more accurate estimation of the parameters using limited real-world data than that learned using a supervised method with a large number of synthetic data. Also, the generalization of the model trained is stronger than that trained using a supervised method with limited real-world data. On the other hand, the unsupervised scheme exploits the unlabeled real-world underwater images that we collect. A second set of images with different degrees of degradation are generated from these images based on the revised IFM [2]. Two sets of parameters can be estimated using both sets of images, respectively. Each set is utilized as the reference for the other set because their IFMs are related.
@@
Figure 3: The proposed semi-supervised learning framework, which comprises a bi-directional supervised learning scheme and an unsupervised learning scheme, used for training our PATS-UIENet.
@@
Due to the ill-posed nature of the revised IFM, however, only $\mathcal{L}_{fwd}$ may be insufficient for ensuring the validity of the estimated parameters. Therefore, we propose a backward-degradation learning module which degrades clear reference images using the estimated parameters to their degraded counterparts $\hat{I}^{c}(x)$. The supervision over the backward-degradation is achieved by minimizing the loss function:
@@
Unsupervised Learning. To leverage unlabeled underwater images, generative networks [30, 19] were used. However, they were difficult to train and produced unstable results. Since unsupervised learning methods [32, 33] are able to directly learn from degraded images without requiring reference images, we are motivated to adopt an unsupervised learning scheme on top of the revised IFM [2].
Given an unlabeled real-world degraded image $I^{c}_{1}(x)$, the revised IFM can be expressed as:
$I^{c}_{1}(x)=J^{c}_{1}(x)t_{D1}^{c}(x)+(1-t_{B1}^{c}(x))A^{c}_{1}.$ (8)
A new degraded image $I^{c}_{2}(x)$ can be purposely derived according to:
$I^{c}_{2}(x)=\alpha I^{c}_{1}(x)+(1-\alpha)(1-t_{B1}^{c}(x))A^{c}_{1},$ (9)
@@
$I^{c}_{2}(x)=J^{c}_{1}(x)(\alpha t_{D1}^{c}(x))+(1-t_{B1}^{c}(x))A^{c}_{1}.$ (10)
Eq. (10) also satisfies the revised IFM but presents a more severe degradation than the $I^{c}_{1}(x)$, generated from the same underlying clear image $J^{c}_{1}(x)$, because $\alpha t_{D1}^{c}(x)<t_{D1}^{c}(x)$ always holds true. Therefore, the parameters $\hat{t}_{D2}^{c}(x)$ and $\hat{t}_{B2}^{c}(x)$ estimated using the PATS-UIENet from image $I^{c}_{2}(x)$ will be close to $\alpha\hat{t}_{D1}^{c}(x)$ and $\hat{t}_{B1}^{c}(x)$, respectively, where $\hat{t}_{D1}^{c}(x)$ and $\hat{t}_{B1}^{c}(x)$ are the parameters estimated from $I^{c}_{1}(x)$. Correspondingly, two loss functions are defined as:
$\mathcal{L}_{T}=||\hat{t}_{D2}^{c}(x)-\alpha\hat{t}_{D1}^{c}(x)||_{2}^{2},$ (11)
$\mathcal{L}_{A\text{-}unsup}=||\hat{t}_{B1}^{c}(x)-\hat{t}_{B2}^{c}(x)||_{2}^{2}.$ (12)
@@ IV-A Baselines
We compared the proposed method with four prior-based approaches which were developed on top of the IFM [1], including DCP [3], UDCP [4], GDCP [5] and DHL [10], four prior-based methods which did not consider the IFM, including Histogram Equalization (HE) [9], Retinex [13], MMLE [14] and HLRP [15], and eight deep learning approaches, including WaterNet [22], UColor [23], FUnIE-GAN [20], USUIR [24], PUIE-Net (MP) [25], PUIE-Net (MC) [25], U-Transformer [21] and URanker [26].
IV-B Data Sets
@@
We conducted a series of experiments using five data sets, including four publicly available real-world underwater data sets, i.e., UIEB [22], SUIM-E [27], RUIE [53] and SQUID [10], and a synthetic data set, namely, EUVP [20]. The UIEB data set originally contains 890 pairs of degraded images, in which each reference image was selected from the results produced by applying 12 algorithms to a degraded image. Ten pairs of UIEB [22] images were discarded because the degraded images in these pairs had been included in the rest pairs. In total, we only used 880 pairs of UIEB images.
The SUIM-E [27] data set comprises 1,525 pairs of degraded underwater images and a testing set of 110 labeled images. Three subsets are comprised of the RUIE [53] data set, including UIQS, UCCS and UHTS, which include 3,630, 300 and 300 unlabeled images, respectively. The SQUID [10] data set comprises 114 underwater images captured using a stereo camera, divided into the Michmoret, Katzaa, Nachsholim and Satil subsets. We utilized 2,185 images from the Underwater Scenes subset of the EUVP [20] data set.
Two different PATS-UIENet models were trained for the real-world and synthetic testing images, referred to as $PATS\text{-}UIENet_{Real}$ and $PATS\text{-}UIENet_{Syn}$, respectively. Regarding both the models, we randomly selected 720 pairs from the 880 pairs of UIEB [22] images and 2,000 pairs from the Underwater Scenes subset of the EUVP [20] data set, respectively, which were used as the training images in the supervised learning scheme. We collected 1,934 unlabeled real-world underwater images, which cover diverse underwater scenes, and used these images in the unsupervised learning scheme for both the models.
@@
Among the rest of the UIEB [22] data set, 80 pairs of images were randomly selected as the validation set, while the remaining pairs were used as the testing set which is referred to as Test-U80. The UIEB data set also consists of 60 challenging images without references, which were used as the second testing set, referred to as Test-C60. A third testing set was obtained from the testing set of the SUIM-E data set [27], which contains 110 pairs of labeled images and is named as Test-S110. For the RUIE [53] data set, the UCCS was utilized as the fourth testing set, denoted as Test-UCCS. Fifty-three SQUID [10] images captured using the camera at the right hand side were randomly selected and were used as the fifth testing set, namely, Test-R53. The remaining 185 images in the Underwater Scenes subset of the EUVP [20] data set were used as the sixth testing set, referred to as Test-Scenes.
IV-C Evaluation Metrics
@@
For the labeled testing data, we used both Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [54] to assess the quality of an enhanced image with regard to the reference image. Regarding the unlabeled testing data, we used the Underwater Image Quality Measure (UIQM) [55] and Underwater Color Image Quality Evaluation (UCIQE) [56] metrics. When the SQUID [10] data set was tested, the Average Reproduction Angular Error ($\bar{\psi}$) [10] was used to evaluate the quality of color restoration. Since transmission maps are related to the scene distance, we used the Pearson Correlation Coefficient (PCC) calculated between the estimated transmission maps and the ground-truth depth map to assess the performance of transmission estimation [10].
IV-D Implementation Details
@@
We implemented the proposed method using Pytorch and conducted experiments on Ubuntu 20.04 with a GeForce RTX 3090 graphics processing unit. During the training process, the images were resized to the resolution of $256\times 256$ pixels. The number of filters in the RCT was set to 16. For the $Enc\_1$ to $Dec\_1$ in the D-stream and B-stream, the numbers of filters were set to 64, 128, 256, 512, 512, 256, 128, 64 and 64 in turn. Regarding the A-stream, the dimension of each token was 384 and the number of attention heads in the MHSA was 6. We trained the PATS-UIENet using the AdamW [57] optimizer. The learning rate and batch size were set to 1e-4 and 12, respectively. Four weighting factors, including $\lambda_{1},\lambda_{2},\lambda_{3}$ and $\lambda_{unsup}$, were set to 0.001, 0.005, 10 and 0.001, respectively. We first trained the PATS-UIENet using the bi-directional supervised scheme for 3,000 iterations as a warm-up stage. Then the semi-supervised scheme was used to train it for 150,000 iterations. When the semi-supervised learning process was carried out, in particular, the unsupervised scheme was performed for 30 iterations after the bi-directional supervised scheme was conducted for every 120 iterations.
V Experimental Results
@@
In this section, we will report the results obtained in the underwater image enhancement, transmission estimation and color restoration experiments and the ablation study.
V-A Underwater Image Enhancement
We evaluated the proposed PATS-UIENet along with 16 baselines using two full-reference quantitative metrics and two non-reference quantitative metrics. A qualitative analysis was also performed. In this subsection, the results obtained in the three experiments will be reported.
V-A1 Full-Reference Quantitative Evaluation
Since the Test-U80, Test-S110, and Test-Scenes testing sets contain reference images, a full-reference quantitative evaluation was conducted on the proposed method and the 16 baselines using the PSNR and SSIM metrics across these testing sets. The results are shown in Table I. As can be seen, our PATS-UIENet achieved the best result in terms of both the metrics no matter which testing set was used. Compared with the prior-based baselines using IFM [1], such as DCP [3], UDCP [4], GDCP [5] and DHL [8], our method showed a significant improvement.
@@ V-A1 Full-Reference Quantitative Evaluation
TABLE I: The full-reference quantitative evaluation of the proposed PATS-UIENet and sixteen baselines on three real-world testing sets. The best and second best results are highlighted in the Red Bold and Blue Italic fonts, respectively.
Method Test-U80 Test-S110 Test-Scenes
@@
DCP 12.48 63.34 14.52 67.38 18.09 68.95
UDCP 9.37 33.62 10.04 34.28 14.76 56.43
GDCP 12.46 67.41 14.67 71.29 14.11 63.50
DHL 14.49 62.81 14.96 69.68 15.16 53.92
HE 16.48 77.77 15.37 74.76 13.61 62.82
Retinex 10.38 61.25 8.29 53.44 9.50 55.56
MMLE 18.56 76.21 17.32 77.39 14.89 62.23
HLRP 13.56 22.76 12.55 29.24 12.12 18.21
WaterNet 17.02 70.27 18.80 75.19 25.84 84.02
UColor 20.93 85.19 20.23 84.23 24.31 81.40
FUnIE-GAN 15.31 59.68 16.89 66.32 25.11 83.16
USUIR 16.92 68.69 17.54 76.58 15.14 66.51
PUIE-Net (MP) 22.26 88.89 21.85 89.19 21.85 73.47
PUIE-Net (MC) 22.05 88.80 21.72 89.20 20.85 73.10
U-Transformer 20.98 72.76 19.92 68.25 24.09 79.98
URanker 21.87 85.75 21.27 87.43 22.46 82.35
PATS-UIENet (Ours) 23.01 89.89 22.48 90.26 27.20 88.30
V-A2 Non-reference Quantitative Evaluation
We used two non-reference metrics, i.e., UIQM [55] and UCIQE [56], to evaluate the performance of our method and sixteen baselines on the Test-U80, Test-S110, Test-Scenes, Test-C60, Test-UCCS and Test-R53 testing sets. The results are shown in Tables II. It can be seen that the performance of our method was lower than the best counterpart but was comparable to most baselines. As demonstrated in [22, 27, 35, 26], the UIQM and UCIQE metrics are not always consistent with the quality of images processed by UIE methods. These findings may explain the relatively lower UIQM and UCIQE values produced by our method.
TABLE II: The non-reference quantitative evaluation of the proposed PATS-UIENet and sixteen baselines on six real-world testing sets. The best and second best results are highlighted in the Red Bold and Blue Italic fonts, respectively.
Method Test-U80 Test-S110 Test-Scenes Test-C60 Test-UCCS Test-R53
[…]
DCP 1.35 0.51 1.36 0.57 1.61 0.57 0.56 0.51 0.03 0.44 -2.05 0.42
UDCP 2.69 0.59 1.23 0.60 2.80 0.60 1.24 0.54 1.83 0.57 0.76 0.59
GDCP 2.17 0.56 2.12 0.64 1.89 0.65 1.49 0.60 1.24 0.56 5.50 0.50
DHL 3.19 0.64 2.35 0.66 3.54 0.66 1.85 0.64 2.98 0.63 0.22 0.64
HE 3.58 0.67 2.97 0.68 4.27 0.67 2.32 0.65 3.93 0.67 1.71 0.64
Retinex 2.33 0.48 0.85 0.45 3.36 0.47 1.21 0.46 2.03 0.47 -2.03 0.51
MMLE 3.89 0.58 2.72 0.60 3.60 0.61 2.77 0.57 4.06 0.57 3.18 0.55
HLRP 3.71 0.63 2.88 0.65 3.85 0.66 1.98 0.65 3.89 0.64 1.98 0.65
WaterNet 2.11 0.50 1.47 0.54 3.67 0.57 1.28 0.51 1.25 0.45 0.09 0.45
UColor 3.44 0.60 2.30 0.63 3.64 0.57 2.08 0.58 3.22 0.55 1.28 0.54
FUnIE-GAN 3.73 0.50 3.59 0.57 4.43 0.58 4.50 0.51 3.47 0.49 0.19 0.42
USUIR 4.10 0.63 2.84 0.65 3.88 0.65 3.12 0.62 4.19 0.60 3.31 0.60
PUIE-Net (MP) 3.42 0.56 2.25 0.59 3.54 0.57 1.82 0.55 4.03 0.52 1.56 0.51
PUIE-Net (MC) 3.31 0.56 2.22 0.59 3.11 0.57 1.78 0.55 3.82 0.52 1.30 0.49
U-Transformer 4.37 0.55 3.68 0.56 4.54 0.55 4.19 0.53 4.19 0.53 2.43 0.53
URanker 3.64 0.57 2.55 0.59 3.68 0.57 2.36 0.56 4.17 0.56 2.25 0.56
PATS-UIENet (Ours) 3.76 0.57 2.47 0.60 3.70 0.57 2.24 0.56 3.58 0.55 2.55 0.55
V-A3 Qualitative Analysis
The visual results generated by the sixteen baselines and our method on six testing sets are shown in Figs. 4 - 9, respectively. It can be seen that some methods with the higher UIQM [55] or UCIQE [56] score did not produce visually pleased results. For example, MMLE [14] increased the contrast but produced images with pale colors (see Fig. 4). FUnIE-GAN [20] could not completely remove the color cast in degraded images (see Fig. 7). HE[9], HLRP[15], and USUIR[24] introduced artifacts that reduced the visual quality (see Fig. 8). In contrast, our method not only effectively removed the color cast from degraded images but also significantly improved the visibility of the images. It is demonstrated that the images that our method enhanced still exhibit the satisfactory visual quality even if the UIQM or UCIQE value that our method generated is lower than that produced by the baselines. This observation is consistent with the findings presented in the literature [22, 27, 35, 26].
Figure 4: The results produced by 16 baselines and our method in terms of a degraded image in the Test-U80 testing set.
Figure 5: The results produced by 16 baselines and our method in terms of a degraded image in the Test-S110 testing set.
Figure 6: The results produced by 16 baselines and our method in terms of a degraded image in the Test-Scenes testing set.
Figure 7: The results produced by 16 baselines and our method in terms of a degraded image in the Test-C60 testing set.
Figure 8: The results produced by 16 baselines and our method in terms of a degraded image in the Test-UCCS testing set.
Figure 9: The results produced by 16 baselines and our method in terms of a degraded image in the Test-R53 testing set.
V-B Transmission Estimation and Color Restoration